# Monotonicity of the ultrafilter number function

### v3

@@ 7. Monotonicity at a singular cardinal and its successor
We had seen that $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})$ always holds if $\kappa$ is regular. By Theorem 4.8, if $\kappa$ is singular then it is possible that $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$. On the other hand, we will show there are some restrictions of the failure of monotonicity at singular and its successor.
We use Shelah’s PCF theory. We present some definitions and basic facts. See Shelah [19], Abraham and Magidor [1], and Eisworth [5] for details.
Fact 7.1 (Shelah).
@@
Now we prove the following proposition, which plays a crucial role in this section.
Proposition 7.3.
Let $\kappa$ be a singular cardinal. Let $S$ be a set and $\mathcal{U}$ an ultrafilter over $S$. If $\mathcal{U}$ is $\kappa^{+}$-decomposable and $\mathord{\mathrm{cf}}(\kappa)$-decomposable, then $\mathcal{U}$ is $\kappa$-decomposable as well.
@@
First we check that $\mathcal{U}$ is $(\kappa,\kappa^{+})$-regular. Since $\mathcal{U}$ is $\kappa^{+}$-decomposable, we can find $f:S\to\kappa^{+}$ such that $f_{*}(\mathcal{U})$ is a uniform ultrafilter over $\kappa^{+}$. By Fact 7.2, $f_{*}(\mathcal{U})$ is $(\kappa,\kappa^{+})$-regular. Fix $\mathcal{F}\subseteq f_{*}(\mathcal{U})$ which witnesses that $f_{*}(\mathcal{U})$ is $(\kappa,\kappa^{+})$-regular. Then it is easy to show that $\{f^{-1}(X)\mid X\in\mathcal{F}\}$ also witnesses that $\mathcal{U}$ is $(\kappa,\kappa^{+})$-regular.
@@
It is sufficient to show that there is a non-$\sigma$-complete uniform ultrafilter $\mathcal{V}$ over $\omega\times\kappa$ with $\chi(\mathcal{V})=\mathfrak{u}(\kappa)$.
By the assumption, there is a measurable cardinal $\leq\kappa$. Hence we have $2^{\omega}\leq\kappa$. Fix a uniform ultrafilter $\mathcal{W}$ over $\omega$. Define the filter $\mathcal{V}$ over $\omega\times\kappa$ by: $X\in\mathcal{V}\iff$ there is $Y\in\mathcal{W}$ and $Z\in\mathcal{U}$ with $Y\times Z\subseteq X$. We have that $\chi(\mathcal{V})\leq\chi(\mathcal{U})\times 2^{\omega}=\chi(\mathcal{U})$. We can check $\mathcal{V}$ is not $\sigma$-complete as follows: $X_{n}=[n,\omega)\times\kappa\in\mathcal{V}$ but $\bigcap_{n}X_{n}=\emptyset$. We have to show that $\mathcal{V}$ is an ultrafilter. Take $X\subseteq\omega\times\kappa$. Consider the set $Y=\{n<\omega\mid\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in X\}\in\mathcal{U}\}$.
Case 1: $Y\in\mathcal{W}$. For $n\in Y$, take $Z_{n}=\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in X\}\in\mathcal{U}$. Since $\mathcal{U}$ is $\sigma$-complete, we have $Z=\bigcap_{n}Z_{n}\in\mathcal{U}$. Then $Y\times Z\subseteq X$, hence $X\in\mathcal{V}$.
Case 2: $Y\notin\mathcal{W}$. Then $Y^{\prime}=\{n<\omega\mid\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in\kappa\setminus X\}\in\mathcal{U}\}$ is in $\mathcal{W}$. For $n\in Y^{\prime}$, let $Z_{n}^{\prime}=\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in\kappa\setminus X\}\in\mathcal{U}$, and $Z^{\prime}=\bigcap_{n}Z_{n}^{\prime}\in\mathcal{U}$. Then $Y^{\prime}\times Z^{\prime}\subseteq(\omega\times\kappa)\setminus X$, hence $(\omega\times\kappa)\setminus X\in\mathcal{V}$. ∎
@@
Suppose to the contrary that $\mathfrak{u}(\kappa)>\mathfrak{u}(\kappa^{+})$. If $\mathcal{U}$ is a uniform ultrafilter over $\kappa^{+}$ with $\chi(\mathcal{U})=\mathfrak{u}(\kappa^{+})$, then $\mathcal{U}$ is $\kappa$-indecomposable but $\kappa^{+}$-decomposable, so $\omega$-indecomposable by Proposition 7.3. In particular $\mathcal{U}$ is $\sigma$-complete. By Lemma 7.4, there is a uniform ultrafilter $\mathcal{V}$ over $\kappa^{+}$ such that $\chi(\mathcal{V})=\mathfrak{u}(\kappa^{+})$ and $\mathcal{V}$ is not $\sigma$-complete. Then $\mathcal{V}$ is $\mathord{\mathrm{cf}}(\kappa)$-decomposable, hence is $\kappa$-decomposable by Proposition 7.3 again, and we have $\mathfrak{u}(\kappa)\leq\chi(\mathcal{V})=\mathfrak{u}(\kappa^{+})$ by Lemma 5.1. This is a contradiction. ∎
We note that, by Theorem 4.8, the countable cofinality assumption of this theorem cannot be eliminated.
@@
Combining Theorem 7.5 with Theorems 4.7 and 6.10 we have:
@@
We may assume $\lambda\geq\mathord{\mathrm{cf}}(\kappa)$. Fix a $\kappa$-indecomposable uniform ultrafilter $\mathcal{U}$ over $\kappa^{+}$, which is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable by Proposition 7.3. $\mathcal{U}$ is $(\kappa,\kappa^{+})$-regular by Fact 7.2, so we can take a family $\{x_{\alpha}\mid\alpha<\kappa^{+}\}$ such that $\left|{x_{\alpha}}\right|<\kappa$ and $\{\alpha\mid i\in x_{\alpha}\}\in\mathcal{U}$ for $i<\kappa^{+}$.
Fix an enumeration $\{B_{i}\mid i<\kappa^{+}\}$ of $\bigcup_{\mu<\kappa}[\mu]^{\leq\lambda}$ such that for every $j<\kappa^{+}$, the set $\{i\mid B_{i}=B_{j}\}$ is cofinal in $\kappa^{+}$. For $\alpha<\kappa^{+}$, let $\mathcal{B}_{\alpha}=\{\bigcup_{i\in c}B_{i}\mid c\in[x_{\alpha}]^{\mathord{\mathrm{cf}}(\kappa)}\}$. Since $\mu^{\lambda}\leq\kappa^{+}$ for every $\mu<\kappa$ and $\left|{x_{\alpha}}\right|<\kappa$, we have $\left|{\mathcal{B}_{\alpha}}\right|\leq\kappa^{+}$. Thus $\left|{\bigcup_{\alpha}\mathcal{B}_{\alpha}}\right|\leq\kappa^{+}$. We prove that $[\kappa]^{\lambda}\subseteq\bigcup_{\alpha}\mathcal{B}_{\alpha}$, then we have $\kappa^{\lambda}\leq\kappa^{+}$ as required.
@@
If $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$, then $\kappa^{+}$ has a $\kappa$-indecomposable uniform ultrafilter, hence $2^{\kappa}=\kappa^{\mathord{\mathrm{cf}}(\kappa)}=\kappa^{+}$ by Proposition 7.8. But then $\mathfrak{u}(\kappa)\leq 2^{\kappa}=\kappa^{+}\leq\mathfrak{u}(\kappa^{+})$, this is a contradiction. ∎
@@
Next we prove that if $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$, then $\mathfrak{u}(\kappa^{+})$ is an upper bound of the ultrafilter numbers at almost all regular cardinals below $\kappa$.
@@
Let $\kappa$ be a singular cardinal, and $A\subseteq\kappa$ be a set of regular cardinals with $\sup(A)=\kappa$ and $\left|{A}\right|=\mathord{\mathrm{cf}}(\kappa)$. Let $U$ be an ultrafilter over $A$. For $f,g\in\prod A$, define $f\leq_{U}g$ by $\{\mu\in A\mid f(\mu)\leq g(\mu)\}\in U$. A set $\mathcal{F}\subseteq\prod A$ is $\leq_{U}$-cofinal if for every $f\in\prod A$, there is $g\in\mathcal{F}$ with $f\leq_{U}g$. Let $\mathord{\mathrm{cf}}(\prod A,\leq_{U})=\min\{\left|{\mathcal{F}}\right|\mid\mathcal{F}$ is $\leq_{U}$-cofinal}}\}}, and $\mathord{\mathrm{PP}}(\kappa)$ the set of all $\mathord{\mathrm{cf}}(\prod A,\leq_{U})$’s where $A\subseteq\kappa$ is a set of regular cardinals with $\sup(A)=\kappa$, $\left|{A}\right|=\mathord{\mathrm{cf}}(\kappa)$, and $U$ an ultrafilter over $A$ which does not contain bounded subsets of $A$. Let $\mathord{\mathrm{pp}}(\kappa)$ be the supremum of $\mathord{\mathrm{PP}}(\kappa)$.
Fact 7.11 (Shelah).
For a singular cardinal $\kappa$, the following are equivalent:
@@
    $\mathord{\mathrm{pp}}(\kappa)=\kappa^{+}$.
@@
    For every increasing sequence $\langle{\kappa_{\xi}\mid\xi<\mathord{\mathrm{cf}}(\kappa)}\rangle$ of regular cardinals with limit $\kappa$, there is a subsequence $\langle{\kappa^{\prime}_{\xi}\mid\xi<\mathord{\mathrm{cf}}(\kappa)}\rangle$ of the sequence with limit $\kappa$ which admits a scale of length $\kappa^{+}$.
Definition 7.12.
For a cardinal $\kappa$, let $\mathrm{ADS}_{\kappa}$ be the following assertion: there is a sequence $\langle{A_{i}\mid i<\kappa^{+}}\rangle$ such that $A_{i}$ is an unbounded subset of $\kappa$ with size $\mathord{\mathrm{cf}}(\kappa)$, and for every $\alpha<\kappa^{+}$ there is $f:\alpha\to\kappa$ such that $(A_{i}\setminus f(i))\cap(A_{j}\setminus f(j))=\emptyset$ for every $i<j<\alpha$.
Fact 7.13 (Shelah).
Let $\kappa$ be a singular cardinal. If $\mathord{\mathrm{pp}}(\kappa)>\kappa^{+}$, then $\mathrm{ADS}_{\kappa}$ holds.
Proposition 7.14.
Let $\kappa$ be a singular cardinal, and suppose $\kappa^{+}$ has a $\kappa$-indecomposable uniform ultrafilter. Then $\mathrm{ADS}_{\kappa}$ fails, in particular $\mathord{\mathrm{pp}}(\kappa)=\kappa^{+}$.
@@
Suppose to the contrary that $\mathrm{ADS}_{\kappa}$ holds. Let $\langle{A_{i}\mid i<\kappa^{+}}\rangle$ be a $\mathrm{ADS}_{\kappa}$-sequence.
Let $\mathcal{U}$ be a $\kappa$-indecomposable uniform ultrafilter over $\kappa^{+}$. By Proposition 7.3, $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable. For $\alpha<\kappa^{+}$, we can find $g_{\alpha}:\alpha+1\to\kappa$ such that the sequence $\langle{A_{i}\setminus f(i)\mid i\leq\alpha}\rangle$ is pairwise disjoint. Fix $\beta<\kappa^{+}$. Since $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable, we can find $\delta_{\beta}<\kappa$ such that $X_{\beta}=\{\alpha<\kappa^{+}\mid g_{\alpha}(\beta)<\delta_{\beta}\}\in\mathcal{U}$. Now there must be some $\delta<\kappa^{+}$ such that $Y=\{\beta<\kappa^{+}\mid\delta_{\beta}=\delta\}$ is unbounded in $\kappa^{+}$. We show that the sequence $\langle{A_{\beta}\setminus\delta\mid\beta\in Z}\rangle$ is pairwise disjoint, but this is impossible since $\left|{Y}\right|=\kappa^{+}$
Pick $\beta,\beta^{\prime}\in Y$ with $\beta<\beta^{\prime}$. We know that $X_{\beta}\cap X_{\beta^{\prime}}\in\mathcal{U}$, so there is a large $\alpha>\beta,\beta^{\prime}$ with $\alpha\in X_{\beta}\cap X_{\beta^{\prime}}$. Then $g_{\alpha}(\beta),g_{\alpha}(\beta^{\prime})<\delta$. Because $A_{\beta}\setminus g_{\alpha}(\beta)$ and $A_{\beta^{\prime}}\setminus g_{\alpha}(\beta^{\prime})$ are disjoint and $g_{\alpha}(\beta),g_{\alpha}(\beta^{\prime})<\delta$, we have that $A_{\beta}\setminus\delta$ and $A_{\beta^{\prime}}\setminus\delta$ are disjoint as well. ∎
Proposition 7.15.
Let $\kappa$ be a singular cardinal, and $\mathcal{U}$ a uniform ultrafilter over $\kappa^{+}$. If $\mathcal{U}$ is $\kappa$-indecomposable, then there is a cardinal $\lambda<\kappa$ such that $\mathcal{U}$ is $\mu$-decomposable for every regular $\mu$ with $\lambda<\mu<\kappa$.
Proof.
Suppose to the contrary that the set $\{\mu<\kappa\mid\mu$ is regular, $\mathcal{U}$ is $\mu$-indecomposable}}\}} is unbounded in $\kappa$.
By Proposition 7.3, we have that $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable. Applying Proposition 7.14 and Facts 7.11, 7.13, we can take an increasing sequence $\langle\kappa_{\xi}\mid\xi<\mathord{\mathrm{cf}}(\kappa)\rangle$ of regular cardinals with limit $\kappa$ such that $\mathcal{U}$ is $\kappa_{\xi}$-indecomposable, and $\prod_{\xi}\kappa_{\xi}$ admits a scale $\langle f_{i}\mid i<\kappa^{+}\rangle$.
For $\xi<\mathord{\mathrm{cf}}(\kappa)$, since $\mathcal{U}$ is $\kappa_{\xi}$-indecomposable, we can find $\delta_{\xi}<\kappa_{\xi}$ such that $\{\alpha<\kappa^{+}\mid f_{\alpha}(\xi)<\delta_{\xi}\}\in\mathcal{U}$. Now take a large $\beta<\kappa^{+}$ such that $\delta_{\xi}<f_{\beta}(\xi)$ for almost all $\xi<\mathord{\mathrm{cf}}(\kappa)$.
For $\alpha<\kappa^{+}$ with $\alpha>\beta$, there is $\zeta_{\alpha}<\mathord{\mathrm{cf}}(\kappa)$ such that $f_{\beta}(\xi)<f_{\alpha}(\xi)$ for all $\xi\geq\zeta_{\alpha}$. Since $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable, there is $\zeta<\mathord{\mathrm{cf}}(\kappa)$ with $\{\alpha<\kappa^{+}\mid\zeta_{\alpha}<\zeta\}\in\mathcal{U}$. Hence we can find $\alpha<\kappa^{+}$ with $\beta<\alpha$ and $\zeta_{\alpha}<\zeta$. By the choice of $\zeta_{\alpha}$, we have $\delta_{\zeta}<f_{\beta}(\zeta)<f_{\alpha}(\zeta)$, this is a contradiction. ∎
The following theorem is immediate from Lemmas 5.1, 5.2 and Proposition 7.15.
@@
Note that we can improve “every regular $\mu$” in this theorem to “every regular or singular with countable cofinality $\mu$” by Theorem 7.5, however cannot to “every cardinal $\mu$”; Suppose $\kappa$ is strongly compact, and let $\mu>\kappa$ be a strong limit singular cardinal with cofinality $\omega_{1}$. Then $\mathrm{Add}(\omega,\mu)$ forces $\mathfrak{u}(\kappa^{+\omega_{1}^{2}+1})<\mathfrak{u}(\kappa^{+\omega_{1}^{2}})$ and $\mathfrak{u}(\kappa^{+\omega_{1}^{2}+1})<\mathfrak{u}(\kappa^{+\omega_{1}\cdot(\alpha+1)})$ for every $\alpha<\omega_{1}$.
@@
Let $\kappa$ be a singular cardinal. Then $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})^{\mathord{\mathrm{cf}}(\kappa)}$.

### v4

@@ 7. Monotonicity at a singular cardinal and its successor
We had seen that $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})$ always holds if $\kappa$ is regular. In contrast with a regular cardinal, if $\kappa$ is singular then it is possible that $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$; By Theorem 4.8, it is consistent that $\mathfrak{u}(\kappa^{+\omega_{1}+1})<\mathfrak{u}(\kappa^{+\omega_{1}})$. Nonetheless, we will show there are some restrictions of the failure of monotonicity at singular and its successor, and prove that if $\kappa$ is singular and $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$, then the cofinality of $\kappa$ must be uncountable and $\kappa$ is never strong limit.
To accomplish that, we will make use of Shelah’s PCF theory. Let us start this section by presenting some basic definitions and facts. See Shelah [23], Abraham and Magidor [1], and Eisworth [7] for details.
Theorem 7.1 (Shelah, see, e.g., Theorem 2.23 and 2.26 in [7]).
@@
Now we prove the following theorem, which plays a crucial role in this section and is interesting in its own right.
Theorem 7.3.
Let $\kappa$ be a singular cardinal and $\mathcal{U}$ an ultrafilter over a set $S$. If $\mathcal{U}$ is $\kappa^{+}$-decomposable and $\mathord{\mathrm{cf}}(\kappa)$-decomposable, then $\mathcal{U}$ is $\kappa$-decomposable as well.
@@
First we check that $\mathcal{U}$ is $(\kappa,\kappa^{+})$-regular. Since $\mathcal{U}$ is $\kappa^{+}$-decomposable, we can find $f:S\to\kappa^{+}$ such that $f_{*}(\mathcal{U})$ is a uniform ultrafilter over $\kappa^{+}$ by Lemma 2.8. By Theorem 7.2, $f_{*}(\mathcal{U})$ is $(\kappa,\kappa^{+})$-regular. Fix $\mathcal{F}\subseteq f_{*}(\mathcal{U})$ which witnesses that $f_{*}(\mathcal{U})$ is $(\kappa,\kappa^{+})$-regular. Then it is easy to show that $\{f^{-1}(X)\mid X\in\mathcal{F}\}$ also witnesses that $\mathcal{U}$ is $(\kappa,\kappa^{+})$-regular.
@@
It is sufficient to show that there is a non-$\sigma$-complete uniform ultrafilter $\mathcal{V}$ over $\omega\times\kappa$ with $\chi(\mathcal{V})\leq\mathfrak{u}(\kappa)$.
Fix a uniform ultrafilter $\mathcal{W}$ over $\omega$. Define the filter $\mathcal{V}$ over $\omega\times\kappa$ by: $X\in\mathcal{V}\iff$ there is $Y\in\mathcal{W}$ and $Z\in\mathcal{U}$ with $Y\times Z\subseteq X$. We have that $\chi(\mathcal{V})\leq\chi(\mathcal{U})\times 2^{\omega}$. By the assumption that there is a $\sigma$-complete uniform ultrafilter over $\kappa$, there is a measurable cardinal $\leq\kappa$. Hence we have $2^{\omega}\leq\kappa\leq\chi(\mathcal{U})$ and so $\chi(\mathcal{V})\leq\chi(\mathcal{U})\times 2^{\omega}=\chi(\mathcal{U})=\mathfrak{u}(\kappa)$. Thus $\chi(\mathcal{V})\leq\mathfrak{u}(\kappa)$. We can check $\mathcal{V}$ is not $\sigma$-complete as follows: $X_{n}=[n,\omega)\times\kappa\in\mathcal{V}$ but $\bigcap_{n}X_{n}=\emptyset$.
Now we have to show that $\mathcal{V}$ is an ultrafilter. Take $X\subseteq\omega\times\kappa$, and we show that $X\in\mathcal{V}$ or $(\omega\times\kappa)\setminus X\in\mathcal{V}$. Consider the set $Y=\{n<\omega\mid\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in X\}\in\mathcal{U}\}$.
Case 1: $Y\in\mathcal{W}$. For $n\in Y$, let $Z_{n}=\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in X\}\in\mathcal{U}$. Since $\mathcal{U}$ is $\sigma$-complete, we have $Z=\bigcap_{n}Z_{n}\in\mathcal{U}$. Then $Y\times Z\subseteq X$, hence $X\in\mathcal{V}$.
Case 2: $Y\notin\mathcal{W}$. Since $\mathcal{U}$ and $\mathcal{W}$ are ultrafilters, the set $Y^{\prime}=\{n<\omega\mid\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in\kappa\setminus X\}\in\mathcal{U}\}$ is in $\mathcal{W}$. For $n\in Y^{\prime}$, let $Z_{n}^{\prime}=\{\alpha<\kappa\mid\langle{n,\alpha}\rangle\in\kappa\setminus X\}\in\mathcal{U}$. Again, since $\mathcal{U}$ is $\sigma$-complete, we have $Z^{\prime}=\bigcap_{n}Z_{n}^{\prime}\in\mathcal{U}$. Then $Y^{\prime}\times Z^{\prime}\subseteq(\omega\times\kappa)\setminus X$, and $(\omega\times\kappa)\setminus X\in\mathcal{V}$. ∎
@@
Suppose to the contrary that $\mathfrak{u}(\kappa)>\mathfrak{u}(\kappa^{+})$. By Lemma 7.4, there is a uniform ultrafilter $\mathcal{U}$ over $\kappa^{+}$ such that $\chi(\mathcal{U})=\mathfrak{u}(\kappa^{+})$ and $\mathcal{U}$ is not $\sigma$-complete. Then $\mathcal{U}$ is $\kappa$-indecomposable by Corollary 5.2. However, since $\mathcal{U}$ is $\kappa^{+}$-decomposable, $\mathcal{U}$ is $\omega$-indecomposable by Theorem 7.3, so $\mathcal{U}$ is $\sigma$-complete. This is a contradiction. ∎
We note that, by Theorem 4.8, the countable cofinality assumption of this corollary cannot be eliminated.
@@
Combining Corollary 7.5 with Theorems 4.7 and 6.12 we have:
@@
We may assume $\lambda\geq\mathord{\mathrm{cf}}(\kappa)$. Fix a $\kappa$-indecomposable uniform ultrafilter $\mathcal{U}$ over $\kappa^{+}$, which is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable by Theorem 7.3. $\mathcal{U}$ is $(\kappa,\kappa^{+})$-regular by Theorem 7.2, so we can take a family $\{x_{\alpha}\mid\alpha<\kappa^{+}\}$ such that $\left|{x_{\alpha}}\right|<\kappa$ and $\{\alpha\mid i\in x_{\alpha}\}\in\mathcal{U}$ for $i<\kappa^{+}$.
Fix an enumeration $\{B_{i}\mid i<\kappa^{+}\}$ of $\bigcup_{\mu<\kappa}[\mu]^{\leq\lambda}$ such that for every $j<\kappa^{+}$, the set $\{i\mid B_{i}=B_{j}\}$ is cofinal in $\kappa^{+}$. For $\alpha<\kappa^{+}$, let $\mathcal{B}_{\alpha}=\{\bigcup_{i\in c}B_{i}\mid c\in[x_{\alpha}]^{\mathord{\mathrm{cf}}(\kappa)}\}$. Since $\mu^{\lambda}\leq\kappa^{+}$ for every $\mu<\kappa$ and $\left|{x_{\alpha}}\right|<\kappa$, we have that $\left|{\mathcal{B}_{\alpha}}\right|\leq\kappa^{+}$. Thus $\left|{\bigcup_{\alpha}\mathcal{B}_{\alpha}}\right|\leq\kappa^{+}$. We prove that $\kappa]^{\lambda}\subseteq\bigcup_{\alpha}\mathcal{B}_{\alpha}$, which complete the proof.
@@
If $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$, then $\kappa^{+}$ has a $\kappa$-indecomposable uniform ultrafilter by Corollary 5.2, hence $2^{\kappa}=\kappa^{\mathord{\mathrm{cf}}(\kappa)}=\kappa^{+}$ by Proposition 7.8. But then $\mathfrak{u}(\kappa)\leq 2^{\kappa}=\kappa^{+}\leq\mathfrak{u}(\kappa^{+})$, this is a contradiction. ∎
@@
Next we prove that if $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$, then $\mathfrak{u}(\kappa^{+})$ is an upper bound of the ultrafilter numbers at almost all regular cardinals below $\kappa$. For this sake, we prove some results about (in)decomposable ultrafilters.
Recall that, for a cardinal $\kappa$ and an ultrafilter $\mathcal{U}$, $\mathcal{U}$ is $\kappa$-descendingly incomplete if there are sets $X_{\alpha}\in\mathcal{U}$ ($\alpha<\kappa$) such that for every $\alpha<\beta<\kappa$, $X_{\alpha}\supseteq X_{\beta}$ holds, and $\bigcap_{\alpha<\kappa}X_{\alpha}=\emptyset$. It is easy to see that every $\kappa$-decomposable ultrafilter is $\kappa$-descendingly incomplete, and the converse also holds if $\kappa$ is regular.
Theorem 7.10 (Kunen and Prikry [17]).
Let $\mathcal{U}$ be an ultrafilter and $\kappa$ a singular cardinal. If $\mathcal{U}$ is $\kappa^{+}$-descendingly incomplete but not $\mathord{\mathrm{cf}}(\kappa)$-descendingly incomplete, then there is a cardinal $\lambda<\kappa$ such that $\mathcal{U}$ is $\mu$-descendingly incomplete for every regular $\mu$ with $\lambda<\mu<\kappa$.
This result lead us to the following notion.
@@
Let $\kappa$ be a limit cardinal and $\mathcal{U}$ an ultrafilter. $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable if there is a cardinal $\lambda<\kappa$ such that $\mathcal{U}$ is $\mu$-decomposable for every regular $\mu$ with $\lambda<\mu<\kappa$.
In terms of (in)decomposability, Kunen and Prikry’s theorem can be restated as: If $\mathcal{U}$ is $\kappa^{+}$-decomposable but $\mathord{\mathrm{cf}}(\kappa)$-indecomposable, then $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable. By Theorem 7.3, we can improve Kunen and Prikry’s theorem as follows. Note that every $\mathord{\mathrm{cf}}(\kappa)$-indecomposable ultrafilter is $\kappa$-indecomposable, and the converse does not hold in general.
Proposition 7.12.
Let $\kappa$ be a singular cardinal. If $\mathcal{U}$ is a $\kappa^{+}$-decomposable but $\kappa$-indecomposable ultrafilter, then $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable. In particular, if $\mathcal{U}$ is $\kappa$-indecomposable uniform ultrafilter over $\kappa^{+}$, then $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable.
Proof.
By Theorem 7.3, we have that $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable. Since $\mathord{\mathrm{cf}}(\kappa)$ is regular, $\mathcal{U}$ is not $\mathord{\mathrm{cf}}(\kappa)$-descendingly incomplete. $\mathcal{U}$ is $\kappa^{+}$-decomposable, hence is $\kappa^{+}$-descendingly complete. Then, by Theorem 7.10, there is a cardinal $\lambda<\kappa$ such that for every regular $\mu$ with $\lambda<\mu<\kappa$, $\mathcal{U}$ is $\mu$-descendingly incomplete, in particular $\mu$-decomposable. ∎
Concerning Kunen and Prikry’s theorem, Lipparini showed the following, which can be seen as a converse of Kunen and Prikry’s theorem.
Theorem 7.13 (Lipparini, Theorem 1 and Corollary 7 in [19]).
Let $\kappa$ be a singular cardinal and $\mathcal{U}$ an ultrafilter.
@@
    If $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable, then $\mathcal{U}$ is $\kappa$-decomposable or $\kappa^{+}$-decomposable.
@@
    If $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable, then $\mathcal{U}$ is $\kappa^{+}$-decomposable if and only if $\mathcal{U}$ is almost $\mathop{<}\kappa$-decomposable.
We have the following variant of this theorem, which summarizes Theorem 7.3 and Proposition 7.12.
Corollary 7.14.
Let $\kappa$ be a singular cardinal and $\mathcal{U}$ an ultrafilter. Then the following are equivalent:
  1. (1)
    $\mathcal{U}$ is $\kappa$-indecomposable but almost $\mathop{<}\kappa$-decomposable.
  2. (2)
    $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable but almost $\mathop{<}\kappa$-decomposable.
  3. (3)
    $\mathcal{U}$ is $\kappa$-indecomposable but $\kappa^{+}$-decomposable.
  4. (4)
    $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable but $\kappa^{+}$-decomposable.
@@
(4) $\Rightarrow$ (3) is trivial since every $\mathord{\mathrm{cf}}(\kappa)$-indecomposable ultrafilter is $\kappa$-indecomposable.
(3) $\Rightarrow$ (2). Suppose (3). Then $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable by Theorem 7.3, and is almost $\mathop{<}\kappa$-decomposable by Proposition 7.12.
(2) $\Rightarrow$ (1) is trivial as the direction (4) $\Rightarrow$ (3).
(1) $\Rightarrow$ (4). Suppose (1). By (1) of Theorem 7.13, $\mathcal{U}$ is $\kappa^{+}$-decomposable, and so $\mathcal{U}$ is $\mathord{\mathrm{cf}}(\kappa)$-indecomposable by Theorem 7.3. ∎
Now the following theorem is immediate from Lemma 5.1, Corollary 5.2 and Proposition 7.12.
@@
Note that we can improve “every regular $\mu$” in this theorem to “every regular or singular with countable cofinality $\mu$” by Corollary 7.5, however cannot to “every cardinal $\mu$”; Suppose $\kappa$ is strongly compact, and let $\mu>\kappa^{+\omega_{1}^{2}}$ be a strong limit singular cardinal with cofinality $\omega_{1}$. Then $\mathrm{Add}(\omega,\mu)$ forces $\mathfrak{u}(\kappa^{+\omega_{1}^{2}+1})<\mathfrak{u}(\kappa^{+\omega_{1}^{2}})$ and $\mathfrak{u}(\kappa^{+\omega_{1}^{2}+1})<\mathfrak{u}(\kappa^{+\omega_{1}\cdot(\alpha+1)})$ for every $\alpha<\omega_{1}$.
@@
Let $\kappa$ be a cardinal. Then $\mathfrak{u}(\kappa)\leq\mathfrak{u}(\kappa^{+})^{\mathord{\mathrm{cf}}(\kappa)}$.